OpenStudy (anonymous):

A model is built with a scale of 1 in. = 100,000 ft. How tall would a 35,600 ft mountain be in the model?

OpenStudy (amonoconnor):

100,000ft = (100,000)(12)in. = 1,200,000in. So... 1:1200000 = x:(35600)(12) 1/1,200,000 = x/(427,200) (427,200)(1/1,200,000) = (x/427,200)(427,200) (427,200/1,200,000) = x x = .356 inches

OpenStudy (amonoconnor):

Does this make sense?

OpenStudy (anonymous):

yes

OpenStudy (amonoconnor):

(Thank you for the medal!)

OpenStudy (anonymous):

welcome lol